# Spontaneous phase locking in a broad-area semiconductor laser

### v1

@@ B.2 Analysis of heterodyne measurements
Figure 8 shows the RF-spectra of the heterodyning signals for $400$ mA pump current, obtained from the measured time traces via Fourier transform (see Appendix A.2), which contain both the beat notes between the lasing modes of the BAL and the reference laser and frequency components of the BAL itself. Hence, the RF-spectrum of the signal from the BAL alone is calculated and subtracted as shown in Fig. 8b. It should be emphasized that the RF-power is subtracted in linear scale, not in logarithmic scale which would mean dividing the RF-spectra. In the subtracted RF-spectrum presented in Fig. 8c, the only strong signals left are those of the beat notes between BAL and reference laser, though some residues of the frequency components stemming from the BAL alone still remain. From Fig. 8c we can deduce the spatial profiles of the lasing modes and the beat note frequencies. Since the beat note frequency $f_{m}$ of mode $m$ with optical frequency is $\nu_{m}$ is $f_{m}=|\nu_{m}-\nu_{r,1}|$, where $\nu_{r,1}$ is the optical frequency of the reference laser, we cannot obtain $\nu_{m}$ from Fig. 8c without additional information.
@@
Figure 9: Determination of the optical frequencies of lasing modes. Comparison of measurements with reference wavelength $\lambda_{r,1}=851.142$ nm (top panels a, b) and $\lambda_{r,2}=851.141$ nm (bottom panels, c, d) at $400$ mA. The optical frequency is given with respect to $\lambda_{r,1}$ in all panels. The spectra in the left column were calculated assuming that the lasing modes have longer wavelengths than the reference laser, and the spectra in the right column assuming they have shorter wavelengths. The green arrows indicate the modes that appear at the same relative optical frequency for both reference wavelengths.
@@
A second measurement with a slightly shifted wavelength (optical frequency $\nu_{r,2}$) of the reference laser allows us to determine the $\nu_{m}$ unambiguously as explained in Fig. 9. When $\nu_{m}<\nu_{r,1}$ ($\nu_{m}>\nu_{r,1}$), the optical frequency of mode $m$ is given by $\nu_{m}=\nu_{r,1}-f_{m}$ ($\nu_{m}=\nu_{r,1}+f_{m}$). Figure 9 shows the resulting spectra for both cases ($\nu_{m}<\nu_{r,1}$ in the left column and $\nu_{m}>\nu_{r,1}$ in the right column) and for both reference wavelengths, where the reference for calculating the relative optical frequency is $\nu_{r,1}$ in all cases. If the hypothesis that $\nu_{m}<\nu_{r,1}$ ($\nu_{m}>\nu_{r,1}$) is correct, then mode $m$ must appear at the same relative frequency in both spectra to the left (right). These modes are marked by the green arrows in each column. This procedure allows to unambiguously determine the mode frequencies $\nu_{m}$, even in the presence of ”degenerate” modes with the same $f_{m}$ but different $\nu_{m}$. Moreover, any residues from the imperfect subtraction of the BAL signals are easily identified as such since they do not appear at the same relative frequency in either column.
@@ C.1 Linewidth determination
where $f_{0}$ is the peak frequency, $\Gamma$ the full width at half maximum (FWHM), and $A$ the amplitude. In the case of spatially resolved data, a fit is made for each lateral position $x$, and the linewidth is averaged over $x$ after removing outliers and unsuitable data points. Data points are removed if (i) the fit algorithm did not converge, (ii) the fitted FWHM is unusually small or large, or (iii) the amplitude is small, which indicates low signal strength and hence unreliable fits. Criterion (ii) means that the linewidth is two time larger or smaller than the median linewidth. Criterion (iii) means that the fitted amplitude $A$ is smaller than the highest fitted amplitude divided by ten.
@@ C.2 The narrow beat note in the RF-spectra
Figure 10: Properties of the narrow beat note. a, Frequency of the narrow beat note in the RF-spectrum and b its linewidth as function of the pump current for five different set temperatures of the laser diode mount. All measurements were made with laser A.
The broad-area laser investigated here features one narrow peak in the RF-spectrum that is related to multiplets of 1st- and 2nd-order modes (see Fig. 1b). Figure 10 shows the frequency and linewidth of this peak for five different temperatures. Both frequency and linewidth have a complicated dependency on the pump current, and the peak frequency does not exhibit the square-root scaling expected for the relaxation oscillation frequency. The linewidth of the beat note is smallest in the current region of $250$ to $650$ mA and increases rapidly for higher pump currents, indicating the loss of phase locking. This is also the regime where the multiplets are observed. Both the peak frequency and linewidth exhibit a number of local extrema and sudden slope changes. These indicate that the nonlinear dynamics creating the multiplets and the phase locking depends sensitively on the pump current, possibly in the context of gain competition between an increasing number of transverse and longitudinal modes.
Both the peak frequency and linewidth depend on the temperature. Figure 10 shows data for five different set temperatures of the laser diode mount ranging from $15^{\circ}$ to $23^{\circ}$, where $21^{\circ}$ is the default temperature used for all other measurements. However, the effect of temperatures changes on both frequency and linewidth is relatively small. Moreover, the positions of extrema and sudden slope changes are almost the same for different temperatures. This demonstrates that the significant changes of the peak properties as function of the pump current are mainly due to changes in gain and mode competition rather than temperature changes induced by increased Joule heating.
Figure 11: Dynamics of laser B. a, RF-spectra of the total laser emission for laser B. b, Peak frequency and c linewidth of the narrow RF-peak. The data for laser A (cyan) and laser B (black) are compared. All measurements are for a set temperature $21^{\circ}$.
@@ C.3 Spacings and linewidths of the multiplet modes
Figure 12: Mode spacings in the multiplets. The NN-spacings (red) are compared to $1/2$ the NNN-spacings (black).
@@
Figure 13: Linewidths of the multiplet modes. a, Linewidths as function of the pump current. b, Mode linewidths organized by multiplet. The size and the color of the circles corresponds to the linewidth. Different multiplets for the same pump current are offset laterally for clarity.
Figure 13a shows a general increase of the linewidths of the multiplet modes with pump current. It should be noted that the two data points with almost zero linewidth at $450$ and $500$ mA are artifacts in cases where the Lorentz fit fails, and the mode frequencies are not reliable either in these cases. Since the linewidth of the beat note stays roughly constant around $10$ MHz in this current regime, the linewidth reduction factor (LRF) increases as well (see Fig. 3e). So increasing the pump current results in both an increase in the mode linewidth and in the strength of phase locking.
Figure 13b shows the linewidths for each multiplet separately, where multiplets for the same pump current are slightly offset for clarity. We systematically observe that every second mode in the multiplets is broader than its neighbors as also shown in Fig. 3b.
@@ D.1 Theory of linewidth and phase fluctuations
In the following we discuss how the linewidths of lasing modes and beat signals on a photodetector are related to the phase fluctuations of the involved lasing modes, and how information on possible correlations due to phase locking can be extracted. Following Ref. [60], we neglect amplitude fluctuations of the lasing modes and only consider their phase fluctuations.
We will use the following notation and relations: we denote with $\tilde{y}(f)=\mathcal{F}[y(t)]$ the Fourier transform $\mathcal{F}$ of a complex-valued function $y(t)$. Then the power spectral density (PSD) of $y$ is
$S_{y}(f)=|\tilde{y}(f)|^{2}=\mathcal{F}[C_{f}(\tau)]$ (3)
@@
Next we consider the beat signal of two modes $m$ and $n$ on a photodetector. The total electric field is
@@
with the same assumptions for the phases of the two modes as discussed above. The photocurrent signal $j$ on the detector is $j(t)=rI(t)$ with $r$ the radiant sensitivity and $I=|E_{t}|^{2}$ the light intensity. The first-order autocorrelation function of the current is $C_{j}(\tau)=r^{2}C_{E_{t}}^{(2)}(\tau)$ where
@@
First we consider the case of uncorrelated modes. This is for example the case in heterodyne measurements where the modes of the BAL and the reference laser are of course uncorrelated. In this case we can use that the sum of two independent Gaussian random variables $z_{1}$ and $z_{2}$ is a Gaussian random variable $z_{3}$ whose average is equal to the sum of the averages of $z_{1}$ and $z_{2}$ and whose variance is equal to the sum of the variances of $z_{1}$ and $z_{2}$, that is $\sigma_{z_{3}}^{2}=\sigma_{z_{1}}^{2}+\sigma_{z_{2}}^{2}$. Thus, the random variable $\delta_{\tau}\phi_{mn}$ is Gaussian and has a variance of $2\pi|\tau|(\Gamma_{m}+\Gamma_{n})$, and applying Eq. (8) yields
@@
is also a Lorentz function, where $\Delta\nu_{mn}=(\omega_{m}-\omega_{n})/(2\pi)$. Hence, the linewidth of a heterodyne beat notes we measure is the linewidth of the BAL mode plus the linewidth of the reference laser. Since the latter is $<1$ MHz while the BAL modes have linwidths of $50$ MHz and more, the contribution from the reference laser can be neglected in practice.
Next we consider the case of modes with correlated phase fluctuations, such as the NNN-mode pairs in the multiplets. In that case the variance of $\delta_{t}\phi_{mn}$ is
$\begin{array}[]{rcl}\mathrm{Var}(\delta_{t}\phi_{mn})&=&\mathrm{Var}(\delta_{\tau}\phi_{m})+\mathrm{Var}(\delta_{\tau}\phi_{n})\\ &&-2\mathrm{Cov}(\delta_{\tau}\phi_{m},\delta_{\tau}\phi_{n})\\ \\ &=&2\pi|\tau|(\Gamma_{m}+\Gamma_{n})\\ &&-4\pi|\tau|\sqrt{\Gamma_{m}\Gamma_{n}}\mathrm{Cov}(z_{m},z_{n})\end{array}$ (16)
@@
where $\rho_{x,y}$ is the Pearson correlation coefficient of two variables $x$ and $y$, and $\rho_{m,n}$ is the phase fluctuation correlation (PFC) of modes $m$ and $n$.
At this point it should be noted that the sum of two correlated random Gaussian variables is not necessarily a Gaussian variable itself. Thus, Eq. (8) is not guaranteed to be applicable and the lineshape of the beat note not necessarily a Lorentz function. The calculation of the term $\langle e^{i\delta_{\tau}\phi_{mn}}\rangle$ requires the evaluation of all moments of $\delta_{\tau}\phi_{mn}$ which include higher-order correlations between $\delta_{\tau}\phi_{m}$ and $\delta_{\tau}\phi_{n}$. Hence, a careful analysis of the lineshape of the beat note of two phase locked modes could yield information about higher-order phase fluctuation correlations, not just the linear correlation $\rho_{m,n}$. In practice, the narrow beat note in the RF-spectrum is well fitted by a Lorentz function, and measuring the lineshape in greater detail to observe possible devations from a Lorentz profile is challenging.
@@
in the limit of equal mode linewidths, $\Gamma_{m}\approx\Gamma_{n}$. Equation (20) is plotted in Fig. 14 and shows that a LRF of $F_{\mathrm{red}}=10$ corresponds to phase fluctuation correlation of $\rho_{m,n}=0.9$.
Finally we consider how the phase fluctuation correlations (PFCs) of two laser modes $m$ and $n$ can be determined from the heterodyne signals, that is, their beat notes with the reference laser $r$. The intensity of the beat notes between mode $m$ and the reference laser is
$I_{m}(t)=2\mathrm{Re}[A_{m}A_{r}\cos(\Delta\omega_{mr}+\phi_{mr})]\,,$ (21)
@@
$P_{m}+i\mathcal{H}[P_{m}]=2A_{m}A_{r}\exp\{i(\Delta\omega_{mr}t+\phi_{mr}\}$ (22)
@@
$\theta_{mr}(t)=\arg[P_{m}+i\mathcal{H}[P_{m}]]=\Delta\omega_{mr}t+\phi_{mr}(t)\,.$ (24)
@@
$\phi_{mr}(t)=\theta_{mr}(t)-\langle\frac{d\theta_{mr}}{dt}\rangle_{t}$ (25)
@@
that we denote as $\delta\phi_{m}$ in the following. In practice, $\tau=20$ ps is the sampling time step of the oscilloscope, and $\delta\phi_{m}$ is dominated by the phase fluctuations of the lasing mode since its linewidth $\Gamma_{m}$ is much larger than $\Gamma_{r}$.
@@
where $\mathcal{E}$ denotes the expectation value. Since the two lasing modes and the reference laser are uncorrelated, we find $\sigma_{\delta\phi_{m,n}}=\sqrt{2\pi|\tau|(\Gamma_{m,n}+\Gamma_{r})}$. The product of the relative phase fluctuations in the nominator is
$\begin{array}[]{rcl}\delta\phi_{m}\delta\phi_{n}&=&2\pi|\tau|\left(\Gamma_{r}z_{r}^{2}+\sqrt{\Gamma_{m}\Gamma_{n}}z_{m}z_{n}\right.\\ &&\left.-\sqrt{\Gamma_{m}\Gamma_{r}}z_{m}z_{r}-\sqrt{\Gamma_{n}\Gamma_{r}}z_{n}z_{r}\right)\,.\end{array}$ (28)
@@
Again we simplify by assuming that $\Gamma_{m}\approx\Gamma_{n}$ to obtain
@@
We conclude this section with a remark why one should consider the correlation between the phase fluctuations $\delta_{\tau}\phi_{m,n}(t)$ and not that between the phases $\phi_{m,n}(t)$ themselves. Following Eq. (7), the phase $\phi_{m}(t)$ is a sum over many realizations of a Gaussian random variable $z_{m}(t)$, in other words, the phase performs a one-dimensional random walk. However, it is well known that two random walks can show high (positive or negative) correlations even if the Gaussian variables generating the two random walks are completely uncorrelated [61, 62]. These correlations are spurious in the sense that there is no mathematical or physical reason attached to them. The spurious correlations persist even for arbitrarily long time series [62]. This phenomenon can be explained as follows: a typical random walk with $M$ steps will have a total displacement proportional to $\sqrt{M}$ either in positive or negative direction. Consequently, two random walks going in the same (opposite) direction will show a high correlation (anti-correlation), whereas low correlations are rarely found because only few random walks have a total displacement close to zero. Hence, for random-walk like time series it is good practice to consider the fluctuations around their mean behavior instead when searching for possible correlations [61].

### v2

@@ B.2 Analysis of heterodyne measurements
Figure 11 shows the RF-spectra of the heterodyning signals for $400$ mA pump current, which are the modulus square of the Fourier transform of the measured time traces [see Eq. (1)]. They contain both the beat notes between the lasing modes of the BAL and the reference laser and frequency components of the BAL dynamics. Hence, the RF-spectrum of the signal from the BAL alone is calculated and subtracted as shown in Fig. 11b. It should be emphasized that (i) the RF-power is subtracted (i.e., not the complex Fourier transform, but its modulus square), and (ii) that the subtraction is performed in linear scale, not in logarithmic scale which would mean dividing the RF-spectra. Please note that no background subtraction was performed when analyzing and plotting the RF-spectra of the BAL alone. This concerns Figs. 1, 2d and 2e, as well as Fig. 14a.
In the subtracted RF-spectrum presented in Fig. 11c, the only strong signals left are those of the beat notes between BAL and reference laser, though some residues of the frequency components stemming from the BAL alone still remain. From Fig. 11c we can deduce the spatial profiles of the lasing modes and the beat note frequencies. Since the beat note frequency $f_{m}$ of mode $m$ with optical frequency is $\nu_{m}$ is $f_{m}=|\nu_{m}-\nu_{r,1}|$, where $\nu_{r,1}$ is the optical frequency of the reference laser during the first measurement, we cannot obtain $\nu_{m}$ from Fig. 11c without additional information.
@@
Figure 12: Determination of the optical frequencies of lasing modes. Comparison of measurements with reference wavelength $\lambda_{r,1}=851.142$ nm (top panels a and b) and $\lambda_{r,2}=851.141$ nm (bottom panels c and d) at $400$ mA. The optical frequency is given with respect to $\lambda_{r,1}$ in all panels. The spectra in the left column were calculated assuming that the lasing modes have longer wavelengths than the reference laser, and the spectra in the right column assuming they have shorter wavelengths. The green arrows indicate the modes that appear at the same relative optical frequency for both reference wavelengths.
@@
A second measurement with a slightly shifted wavelength (optical frequency $\nu_{r,2}$) of the reference laser allows us to determine the $\nu_{m}$ unambiguously as explained in Fig. 12. When $\nu_{m}<\nu_{r,1}$ ($\nu_{m}>\nu_{r,1}$), the optical frequency of mode $m$ is given by $\nu_{m}=\nu_{r,1}-f_{m}$ ($\nu_{m}=\nu_{r,1}+f_{m}$). Figure 12 shows the resulting spectra for both cases ($\nu_{m}<\nu_{r,1}$ in the left column and $\nu_{m}>\nu_{r,1}$ in the right column) and for both reference wavelengths, where the reference for calculating the relative optical frequency is $\nu_{r,1}$ in all cases. If the hypothesis that $\nu_{m}<\nu_{r,1}$ ($\nu_{m}>\nu_{r,1}$) is correct, then mode $m$ must appear at the same relative frequency in both spectra in the left (right) column. These modes are marked by the green arrows in each column. This procedure allows to unambiguously determine the mode frequencies $\nu_{m}$, even in the presence of ”degenerate” modes with the same $f_{m}$ but different $\nu_{m}$. Moreover, any residues from the imperfect subtraction of the BAL signals are easily identified as such since they do not appear at the same relative frequency in either column.
@@ C.1 Linewidth determination
where $f_{0}$ is the peak frequency, $\Gamma$ the full width at half maximum (FWHM), and $A$ the amplitude. In the case of spatially resolved data, a fit is made for each lateral position $x$, and the linewidth is averaged over $x$ after removing outliers and unsuitable data points. Data points are removed if (i) the fit algorithm does not converge, (ii) the fitted FWHM is unusually small or large, or (iii) the amplitude is small, which indicates low signal strength and hence unreliable fits. Criterion (ii) means that the linewidth is two time larger or smaller than the median linewidth. Criterion (iii) means that the fitted amplitude $A$ is smaller than the highest fitted amplitude divided by ten.
@@ C.2 The narrow beat note in the RF-spectra
Figure 13: Properties of the narrow beat note. a, Frequency of the narrow beat note in the RF-spectrum for five different set temperatures of the laser diode mount. The inset shows the square of the frequency as function of the pump current. b, Linewidth of the beat note as function of the pump current. All measurements were made with laser A.
The broad-area laser (BAL) investigated here features one particularly narrow peak in the RF-spectrum that is related to multiplets of 1st- and 2nd-order modes (see Fig. 1b). Figure 13 shows the frequency and linewidth of this peak as function of the pump current. Both frequency and linewidth have a complicated dependency on the pump current, and the peak frequency does not exhibit the square-root scaling expected for the relaxation oscillation frequency (see inset of Fig. 13a). The linewidth of the beat note is smallest in the current region of $250$ to $650$ mA, which is the regime where the multiplets are observed. It increases rapidly for higher pump currents, indicating the loss of phase locking. Both the peak frequency and linewidth exhibit a number of local extrema and sudden slope changes. These indicate that the nonlinear dynamics creating the multiplets and the phase locking depends sensitively on the pump current, possibly in the context of gain competition between an increasing number of transverse and longitudinal modes.
Figure 14: Dynamics of laser B. a, RF-spectra of the total laser emission for laser B. b, Peak frequency and c, linewidth of the narrow RF-peak. The data for laser A (cyan) and laser B (black) are compared. All measurements are made at a set temperature $21^{\circ}$.
Both the peak frequency and linewidth depend on the temperature. Figure 13 shows data for five different set temperatures of the laser diode mount ranging from $15^{\circ}$ to $23^{\circ}$, where $21^{\circ}$ is the default temperature used for all other measurements. However, the effect of temperatures changes on both frequency and linewidth is relatively small. Moreover, the positions of extrema and sudden slope changes are almost the same for different temperatures. This indicates that the significant changes of the peak properties as function of the pump current are mainly due to changes in gain and mode competition rather than temperature changes induced by increased Joule heating.
@@ C.3 Spacings and linewidths of the multiplet modes
Figure 15: Mode spacings in the multiplets. The NN-spacings (red) are compared to the NNN-spacings divided by two (black). Data are presented as mean +/- standard deviation for $141$ transverse positions.
@@
Figure 16: Linewidths of the multiplet modes. a, Linewidths as function of the pump current. b, Mode linewidths organized by multiplet. The size and the color of the circles correspond to the linewidth. Different multiplets for the same pump current are offset horizontally for clarity.
Figure 16a shows a general increase of the linewidths of the multiplet modes with pump current. It should be noted that the two data points with almost zero linewidth at $450$ and $500$ mA are artifacts in cases where the Lorentz fit fails, and the mode frequencies are not reliable either in these cases. Since the linewidth of the beat note stays roughly constant around $10$ MHz in this current regime, the linewidth reduction factor (LRF) increases as well (see Fig. 3e). So, increasing the pump current results in both an increase in the mode linewidth and in the strength of phase locking.
Figure 16b shows the linewidths for each multiplet separately, where multiplets for the same pump current are horizontally offset for clarity. We systematically observe that every second mode in the multiplets is broader than its neighbors as also shown in Fig. 3b.
@@ D.1 Theory of linewidth and phase fluctuations
In the following we discuss how the linewidths of lasing modes and beat notes on a photodetector are related to the phase fluctuations of the involved lasing modes, and how information on correlations due to phase locking can be extracted. Following Ref. [65], we neglect amplitude fluctuations of the lasing modes and consider only their phase fluctuations.
We will use the following notation and relations: we denote the Fourier transform $\mathcal{F}$ of a complex-valued function $y(t)$ with $\tilde{y}(f)=\mathcal{F}[y(t)]$. Then the power spectral density (PSD) of $y$ is
$S_{y}(f)=|\tilde{y}(f)|^{2}=\mathcal{F}[C_{y}(\tau)]$ (4)
@@
Next, we consider the beat signal of two modes $m$ and $n$ on a photodetector. The total electric field is
@@
with the same assumptions for the phases of the two modes as discussed above. The photocurrent signal $j$ on the detector is $j(t)=rP(t)$ with $r$ the radiant sensitivity and $P=|E_{t}|^{2}$ the optical power. The first-order autocorrelation function of the current is $C_{j}(\tau)=r^{2}C_{E_{t}}^{(2)}(\tau)$ where
@@
First, we consider the case of uncorrelated modes. This is for example the case in heterodyne measurements where the modes of the BAL and the reference laser are of course uncorrelated. In this case we can use that the sum of two independent Gaussian random variables $z_{1}$ and $z_{2}$ is a Gaussian random variable $z_{3}$ whose average is equal to the sum of the averages of $z_{1}$ and $z_{2}$ and whose variance is equal to the sum of the variances of $z_{1}$ and $z_{2}$, that is $\sigma_{z_{3}}^{2}=\sigma_{z_{1}}^{2}+\sigma_{z_{2}}^{2}$. Thus, the random variable $\delta_{\tau}\phi_{mn}$ is Gaussian and has a variance of $2\pi|\tau|(\Gamma_{m}+\Gamma_{n})$, and applying Eq. (9) yields
@@
is also a Lorentz function, where $\Delta\nu_{mn}=(\omega_{m}-\omega_{n})/(2\pi)$. Hence the linewidth of a heterodyne beat note that we measure is the linewidth of the BAL mode plus the linewidth of the reference laser. Since the latter is $<1$ MHz while the BAL modes have typically linewidths of $50$ MHz and more, the contribution from the reference laser can be neglected in practice.
Next, we consider the case of modes with correlated phase fluctuations, such as the NNN-mode pairs in the multiplets which are phase-locked. In that case the variance of $\delta_{\tau}\phi_{mn}$ is
$\begin{array}[]{rcl}\mathrm{Var}(\delta_{\tau}\phi_{mn})&=&\mathrm{Var}(\delta_{\tau}\phi_{m})+\mathrm{Var}(\delta_{\tau}\phi_{n})\\ &&-2\mathrm{Cov}(\delta_{\tau}\phi_{m},\delta_{\tau}\phi_{n})\\ \\ &=&2\pi|\tau|(\Gamma_{m}+\Gamma_{n})\\ &&-4\pi|\tau|\sqrt{\Gamma_{m}\Gamma_{n}}\mathrm{Cov}(z_{m},z_{n})\end{array}$ (17)
@@
where $\rho_{x,y}$ is the Pearson correlation coefficient of two functions $x$ and $y$, and $\rho_{m,n}$ is the phase fluctuation correlation (PFC) of modes $m$ and $n$.
At this point it should be noted that the sum of two correlated random Gaussian variables is not necessarily a Gaussian variable itself. Thus, Eq. (9) is not guaranteed to be applicable and the lineshape of the beat note not necessarily a Lorentz function. The calculation of the term $\langle e^{i\delta_{\tau}\phi_{mn}}\rangle$ requires the evaluation of all moments of $\delta_{\tau}\phi_{mn}$ which include higher-order correlations between $\delta_{\tau}\phi_{m}$ and $\delta_{\tau}\phi_{n}$. Hence, a careful analysis of the lineshape of the beat note of two phase-locked modes could yield information about higher-order phase fluctuation correlations, not just the linear correlation $\rho_{m,n}$. In practice, the narrow beat note in the RF-spectrum is well fitted by a Lorentz function, and measuring the lineshape in greater detail to observe possible deviations from a Lorentz profile is challenging.
@@
in the limit of equal mode linewidths, $\Gamma_{m}\approx\Gamma_{n}$. Equation (21) is plotted in Fig. 17 and shows that a LRF of $F_{\mathrm{red}}=10$ corresponds to a phase fluctuation correlation of $\rho_{m,n}=0.9$.
Finally, we consider how the phase fluctuation correlations of two laser modes $m$ and $n$ can be determined from the heterodyne signals, that is, their beat notes with the reference laser $r$. The power of the beat notes between mode $m$ and the reference laser is
$P_{mr}(t)=2\mathrm{Re}[A_{m}A_{r}\cos(\Delta\omega_{mr}t+\phi_{mr})]\,,$ (22)
@@
$P_{mr}+i\mathcal{H}[P_{mr}]=2A_{m}A_{r}\exp\{i(\Delta\omega_{mr}t+\phi_{mr})\}$ (23)
@@
$\theta_{mr}(t)=\arg\{P_{mr}+i\mathcal{H}[P_{mr}]\}=\Delta\omega_{mr}t+\phi_{mr}(t)\,.$ (25)
@@
$\phi_{mr}(t)=\theta_{mr}(t)-t\left\langle\frac{d\theta_{mr}}{dt}\right\rangle_{t}$ (26)
@@
that we denote as $\delta\phi_{m}$ in the following. In practice, $\tau=20$ ps is the sampling step of the oscilloscope, and $\delta\phi_{m}$ is dominated by the phase fluctuations of the lasing mode since its linewidth $\Gamma_{m}$ is much larger than $\Gamma_{r}$.
@@
where $\mathcal{E}(y)$ denotes the expectation value and $\sigma_{y}$ the standard deviation of a function or variable $y$. Since the two lasing modes are not correlated with the reference laser, we find $\sigma_{\delta\phi_{m,n}}=\sqrt{2\pi|\tau|(\Gamma_{m,n}+\Gamma_{r})}$. The product of the relative phase fluctuations in the nominator is
$\begin{array}[]{rcl}\delta\phi_{m}\delta\phi_{n}&=&2\pi|\tau|\left(\Gamma_{r}z_{r}^{2}+\sqrt{\Gamma_{m}\Gamma_{n}}z_{m}z_{n}\right.\\ \\ &&\left.-\sqrt{\Gamma_{m}\Gamma_{r}}z_{m}z_{r}-\sqrt{\Gamma_{n}\Gamma_{r}}z_{n}z_{r}\right)\,.\end{array}$ (29)
@@
Again, we simplify by assuming that $\Gamma_{m}\approx\Gamma_{n}$ to obtain
@@
We conclude this section with a remark why one should consider the correlations between the phase fluctuations $\delta_{\tau}\phi_{m,n}(t)$ and not that between the phases $\phi_{m,n}(t)$ themselves. Following Eq. (8), the phase $\phi_{m}(t)$ is a sum over many uncorrelated realizations of a Gaussian random variable $z_{m}(t)$, in other words, the phase performs a one-dimensional random walk. However, it is well known that two random walks can show high (positive or negative) correlations even if the Gaussian variables generating the two random walks are completely uncorrelated [66, 67]. These correlations are spurious in the sense that there is no mathematical or physical meaning attached to them. The spurious correlations persist even for arbitrarily long time series [67]. This phenomenon can be explained as follows: a typical random walk with $M$ steps will have a total displacement proportional to $\sqrt{M}$ either in positive or negative direction. Consequently, two random walks going in the same (opposite) direction will show a significant correlation (anti-correlation), whereas low correlations are rarely found because only few random walks have a total displacement close to zero. Hence, for random-walk like time series it is good practice to consider the fluctuations around their mean behavior instead when searching for possible correlations [66].